1.0 g sample of a diatomic gas occupies a volume of 350 mL at 0.0*C and 2.0 atm. Find the mass of

one atom of the gas.

A. 2.7 X 10^23g

B. 2.7 X 10^20g

C. 1.4 X 10^23g

D. 1.4 X 10^20g
Answer #1

Using ideal gas equation,

n= PV/RT

T = 273 K

n = 2*0.35/0.0821*273 = 0.0312 mol

Mass/molar mass = 0.0312 mol

Mass = 1 g

Molar mass = 32.05

This is the mass of one mole of molecules. Mass of 1 mole of atoms of the diatomic gas = 32.05/2 = 16.025 g

Mass of 1 atom = 16.025/6.022*10^23 = 2.7*10^-23

Answer: 2.7*10^-23 g